Suppose four 30g ice are added to a glass containing 200cm3 of juice at 20degree Celsius. The juice has a density of 1gcm3, specific heat of 4.184J/g-K and heat of fusion of 333J/g. Determine whether or not the ice will melt
1
Expert's answer
2020-04-13T02:34:59-0400

m(juice) = V*"{\rho}" = 200*1 = 200 g

The amount of heat that juice can give off is 200*4.184*(20-0) = 16736 J

The amount of heat that ice must absorb to melt (assuming that temperature of ice before adding is 0 0C) is 30*333 = 9990 J

16736>9990, ice will melt

(we neglect the environmental influences (absorbing heat from air or glass))